-2 - 3[2 - 4(x + 1)] = -3 - 5[2 - (x - 5)]=
choose from
a. {-6}
b. {-11/2}
c. {-11/7}
d. {5}

Answer by rfer(16322)

-2-3[2-4x-4]=-3-5[2-x+5]
-2+6+12x=-3-10+5x-25
4+12x=-38+5x
7x=-42
x=-42/7
x=-6
(A)